# Estimating the size of a set using cascading exclusion

### v1

@@ 1.2 An abstract setting
The limits in Theorem 1.1 are developed to prove Lemma 2.2, which implies by Theorem 2.1 that $\mathbb{E}[(n_{0}-\widehat{n}_{0})^{2}]\leq\frac{5}{n-2}$. Section 2.2 gives the promised interpolation between the ‘unknown $k$’ birthday problem and the German tank problem in the presence of a partial order on $S$. Section 2.3 treats the continuous problem of estimating the volume of a convex set. It gives explicit finite sample accuracy bounds that are useful for any dimension $d$ provided that roughly $n\gg d$. Section 2.5 treats a problem suggested by David Aldous. There, $|S|$ is a well defined corpus — e.g., a library of DNA sequences or a collection of melodies. One wants to test if a new entry is suspiciously close to any element of the target collection. The setup uses a distance on $S$ and gives a universally valid test. Section 2.6 treats a general prediction problem where one observes $\{(X_{i},Y_{i})\}_{i=1,\ldots,n}$, builds a prediction interval for $Y$ given $X$ using some given procedure, and then estimates the coverage probability of the prediction interval using a leave-one-out estimator.
@@
All sections contain worked examples with available code (see github). All contain reviews of the literature and can be read now for further motivation.
@@ 2.2.1 Small illustrations in low dimension
For the example in Figure 1, there are 15 extreme points and the convex hull has an area of $7.266$, which gives $\widehat{\mathrm{vol}(K)}=8.55$.
@@ 2.3 Interpolating between birthdays and German tanks using posets
In the introduction, two variants of the problem ‘Estimate $|S|$’ are discussed. In the first (birthday variant), no structure on $S$ is assumed and a sample size $\gg\sqrt{|S|}$ is required. In the second, $S$ is given with a linear order as $S=\{1,2,\ldots,|S|\}$ and a sample of any growing size suffices. This section interpolates between these when $S$ is partially ordered. The ideas are ‘easy’ but since ‘poset theory’ is not standard fare, we proceed slowly. Richard Stanley’s [48, Chapter 3] is a standard reference to partially ordered sets. William Trotter’s Combinatorics and Partially Ordered Sets: Dimension Theory [50] is also useful.
@@
Up-sets are a basic construct of poset theory where they are called ‘order ideals’ (see Stanley [48, Section 3.4]). A simple example of an up-set comes from using the opposite order on $S=\{1,\ldots,N\}$ and the product order on $S\times S$; the up-sets are partitions $\{(i,j),1\leq i\leq I,1\leq j\leq\lambda_{i}\mbox{ with }\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{I}\}$.
@@
What we have called the ‘German tank problem’ was studied earlier by Harold Jeffreys as the tramcar problem [33, section 4.8], for which he used a Bayesian method. Indeed, all problems stated here can be given a Bayesian treatment. We plan to illustrate this in forthcoming work.
@@
Let $S$ be the set of nodes of the infinite rooted labeled tree. We will say that $x\preceq y$ if either $y=x$ or $y$ is an ancestor of $x$. Let $T$ be a finite up-set, which in this case means that $T$ is a finite subtree of $S$ that contains the root. Let $X_{1},\ldots,X_{n}$ be an i.i.d. uniform random sample from $T$. Then $\uparrow\!\{X_{1},\ldots,X_{n}\}$ is the subtree of $S$ consisting of all nodes that are either in the sample or has a descendant in the sample, and $N_{n}$ is the number of sample points which have neither any duplicates nor any descendants in the sample. Then Corollary 2.7 says that
@@
is a good estimator of $|T|$ when $n$ is large and $N_{n}$ turns out to be $\gg\sqrt{n}$. It is not clear whether there is a simple way to see this, especially when nothing is known about the structure of the subtree $T$. Stanley [48, Exercise 3.74] shows that the poset of trees is precisely the set of binary stopping rules. This seems to be worth developing.
We have not seen previous literature on sampling in the presence of a partial order. However, there is extensive literature on computing for posets. For example, a basic theorem of Dilworth says that a finite poset can be covered by a disjoint union of $n$ chains if and only if the size of the largest antichain is $n$. The literature on efficient computation of $n$ and such covers can be found in [2]. One can think about ‘how to estimate the size of a poset’ given a decomposition into disjoint chains as alternatives.
@@ 2.4 Convex subsets of posets
A subset $T\subseteq S$ is called convex if $\operatorname{conv}(T)=T$. An interval $\left[a,b\right]=\{c:a\preceq c\preceq b\}$ is convex. Referring to Example 2.9, the convex subsets of $S\times S$ are the skew-partitions Stanley [48, section 1.10].
@@
Let $S$ be the set of nodes of the infinite rooted labeled tree. As in Example 2.10, will say that $x\preceq y$ if either $y=x$ or $y$ is an ancestor of $x$. Let $T$ be a finite convex subset of $S$, which in this case just means that $T$ is a finite subforest (i.e., union of subtrees) of $S$, with the property that whenever $x,y\in T$ and $x$ is a descendant of $y$, all points in the path from $y$ to $x$ are also in $T$. A key difference with Example 2.10 is that $T$ is no longer required to contain the root. Let $X_{1},\ldots,X_{n}$ be an i.i.d. uniform random sample from $T$. Then $\operatorname{conv}(X_{1},\ldots,X_{n})$ is the subforest of $S$ consisting of all nodes that are either in the sample or has both a descendant and an ancestor in the sample, and $N_{n}$ is the number of sample points which do not have duplicates in the sample and are not sandwiched between two other elements in the sample (that is, an ancestor and a descendant). See Figure 6 for an illustration. Then Corollary 2.12 says that
@@ 2.5.3 Statistical discussion (real world cases)
As an example, that is developed with available R code in the supplementary material, we choose $200$ bacterial DNA sequences from the SILVA [41] database used in DADA2 [5]. We align them and take subsequences of length $450$. We repeat the simulation experiment described above using a reference population of $200$ sequences, pick samples of size $40$ and compute the nearest-neighbor distances using the standard two-parameter Kimura distance [28, Section 10.4], we also compute the $160$ sample-to-population distances as before. This is repeated $500$ times, and we show the two histograms in Figure 9.
@@ Acknowledgements
We thank David Aldous, Eugenio Regazzini, Sandy Zabell and Richard Stanley for their help. PD was supported by NSF grant DMS-1954042. SC was supported in part by NSF grants DMS-2113242 and DMS-2153654.

### v2

@@ 1.2 An abstract setting
The limits in Theorem 1.1 are developed to prove Lemma 2.2, which implies by Theorem 2.1 that $\mathbb{E}[(n_{0}-\widehat{n}_{0})^{2}]\leq\frac{5}{n-2}$. Section 2.2 gives the promised interpolation between the ‘unknown $k$’ birthday problem and the German tank problem in the presence of a partial order on $S$. Section 2.3 treats the continuous problem of estimating the volume of a convex set. It gives explicit finite sample accuracy bounds that are useful for any dimension $d$ provided that roughly $n\gg d$. Section 2.5 treats a problem suggested by David Aldous. There, $|S|$ is a well defined corpus — e.g., a library of DNA sequences or a collection of melodies. One wants to test if a new entry is suspiciously close to any element of the target collection. The setup uses a distance on $S$ and gives a universally valid test. Section 2.6 treats a general prediction problem where one observes $\{(X_{i},Y_{i})\}_{i=1,\ldots,n}$, builds a prediction interval for $Y$ given $X$ using some given procedure, and then estimates the coverage probability of the prediction interval using a leave-one-out estimator. Section 2.7 draws connection to an old theorem of Devroye and Wagner on estimating misclassification rates.
@@
All sections contain worked examples with available code (see github repository here: https://spholmes.github.io/SetSize/). All contain reviews of the literature and can be read now for further motivation.
@@ 2.2.1 Small illustrations in low dimension
For the example in Figure 1, there are 15 extreme points and the convex hull has an area of $7.266$, which gives $\widehat{\mathrm{vol}(K)}=8.55$, where $\mathrm{vol}(K)=8$.
@@ 2.3 Interpolating between birthdays and German tanks using posets
In the introduction, two variants of the problem ‘Estimate $|S|$’ are discussed. In the first (birthday variant), no structure on $S$ is assumed and a sample size $\gg\sqrt{|S|}$ is required. In the second, $S$ is given with a linear order as $S=\{1,2,\ldots,|S|\}$ and a sample of any growing size suffices. This section interpolates between these when $S$ is partially ordered. The ideas are ‘easy’ but since ‘poset theory’ is not standard fare, we proceed slowly. Richard Stanley’s [54, Chapter 3] is a standard reference to partially ordered sets. William Trotter’s Combinatorics and Partially Ordered Sets: Dimension Theory [56] is also useful.
@@
Up-sets are a basic construct of poset theory where they are called ‘order ideals’ (see Stanley [54, Section 3.4]). A simple example of an up-set comes from using the opposite order on $S=\{1,\ldots,N\}$ and the product order on $S\times S$; the up-sets are partitions $\{(i,j),1\leq i\leq I,1\leq j\leq\lambda_{i}\mbox{ with }\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{I}\}$.
@@
What we have called the ‘German tank problem’ was studied earlier by Harold Jeffreys as the tramcar problem [37, section 4.8], for which he used a Bayesian method. Indeed, all problems stated here can be given a Bayesian treatment. We plan to illustrate this in forthcoming work.
A most useful history of the problem of estimating $N$ based on a uniform sample size $n$ appears in a paper of Spencer and Langley [53]. Their focus is on the statistician R. C. Geary’s work on the problem, but along the way they give a host of earlier references, tracing the problem back to C. S. Peirce and Laplace [53, p. 286, footnotes 2,3]. Geary suggested the interesting estimate $\widehat{N}=2^{\frac{1}{n}}\max_{1\leq i\leq n}x_{i}$ and showed that it was ‘closer’ to $N$ than the MLE in the sense of Pitman.
@@
Let $S$ be the set of nodes of an infinite rooted labeled tree. We will say that $x\preceq y$ if either $y=x$ or $y$ is an ancestor of $x$. Let $T$ be a finite up-set, which in this case means that $T$ is a finite subtree of $S$ that contains the root. Let $X_{1},\ldots,X_{n}$ be an i.i.d. uniform random sample from $T$. Then $\uparrow\!\{X_{1},\ldots,X_{n}\}$ is the subtree of $S$ consisting of all nodes that are either in the sample or has a descendant in the sample, and $N_{n}$ is the number of sample points which have neither any duplicates nor any descendants in the sample. Then Corollary 2.7 says that
@@
is a good estimator of $|T|$ when $n$ is large and $N_{n}$ turns out to be $\gg\sqrt{n}$. It is not clear whether there is a simple way to see this, especially when nothing is known about the structure of the subtree $T$. Stanley [54, Exercise 3.74] shows that the poset of trees is precisely the set of binary stopping rules. This seems to be worth developing.
We have not seen previous literature on sampling in the presence of a partial order. Indeed since the poset itself is an up-set, theorem 2.6 gives a novel estimate for the size of a poset. A worthwhile potential application is Dedekind’s problem of estimating the number of order ideals in the Boolean lattice, see [31] for references and first efforts.
However, there is extensive literature on computing for posets. For example, a basic theorem of Dilworth says that a finite poset can be covered by a disjoint union of $n$ chains if and only if the size of the largest antichain is $n$. The literature on efficient computation of $n$ and such covers can be found in [3]. One can think about ‘how to estimate the size of a poset’ given a decomposition into disjoint chains as alternatives.
@@ 2.4 Convex subsets of posets
A subset $T\subseteq S$ is called convex if $\operatorname{conv}(T)=T$. An interval $\left[a,b\right]=\{c:a\preceq c\preceq b\}$ is convex. Referring to Example 2.9, the convex subsets of $S\times S$ are the skew-partitions Stanley [54, section 1.10].
@@
Let $S$ be the set of nodes of an infinite rooted labeled tree. As in Example 2.10, will say that $x\preceq y$ if either $y=x$ or $y$ is an ancestor of $x$. Let $T$ be a finite convex subset of $S$, which in this case just means that $T$ is a finite subforest (i.e., union of subtrees) of $S$, with the property that whenever $x,y\in T$ and $x$ is a descendant of $y$, all points in the path from $y$ to $x$ are also in $T$. A key difference with Example 2.10 is that $T$ is no longer required to contain the root. Let $X_{1},\ldots,X_{n}$ be an i.i.d. uniform random sample from $T$. Then $\operatorname{conv}(X_{1},\ldots,X_{n})$ is the subforest of $S$ consisting of all nodes that are either in the sample or has both a descendant and an ancestor in the sample, and $N_{n}$ is the number of sample points which do not have duplicates in the sample and are not sandwiched between two other elements in the sample (that is, an ancestor and a descendant). See Figure 6 for an illustration. Then Corollary 2.12 says that
@@ 2.5.3 Statistical discussion (real world cases)
As an example, that is developed with available R code in the supplementary material, we choose $200$ bacterial DNA sequences from the SILVA [45] database used in DADA2 [6]. We align them and take subsequences of length $450$. We repeat the simulation experiment described above using a reference population of $200$ sequences, pick samples of size $40$ and compute the nearest-neighbor distances using the standard two-parameter Kimura distance [32, Section 10.4], we also compute the $160$ sample-to-population distances as before. This is repeated $500$ times, and we show the two histograms in Figure 9.
@@ 2.6.3 Literature review
2.7 Connection to a theorem of Devroye and Wagner
It was pointed out to us by Louigi Addario-Berry that our main result bears a resemblance with a theorem proved by Luc Devroye in his Ph.D. thesis [12, Theorem 6.1], based on a prior technical report with his advisor Terry Wagner (see also [47]). An updated version of this theorem appears in the book by Devroye, Györfi, and Lugosi [13, Theorem 24.2]. We now discuss this result and its connection to our Theorem 1.1.
The setting of the Devroye–Wagner theorem is as follows. The data consists of an i.i.d. sequence $D_{n}$ of pairs $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$, where the $X_{i}$’s take value in some arbitrary space and the $Y_{i}$’s are 0-11 valued. The goal is to construct a classifier $g_{n}$ based on this data, which, given a new data point $X_{n+1}$, will output $g_{n}(X_{n+1},D_{n})\in\{0,1\}$ as the predicted value of $Y_{n+1}$. We assume that the classifier is symmetric meaning that $g_{n}(x,D_{n}^{\prime})=g_{n}(x,D_{n})$, where $D_{n}^{\prime}$ is any permutation of $D_{n}$.
Define $L_{n}$ to be the misclassification rate of $g_{n}$ conditional on the data $D_{n}$; that is,
$L_{n}=\mathbb{P}(g_{n}(X_{n+1},D_{n})\neq Y_{n+1}|D_{n}).$
Suppose we have another symmetric classifier $g_{n-1}$ for samples of size n1n-1, and suppose we estimate $L_{n}$ using a leave-one-out procedure based on $g_{n}$:
$\widehat{L}_{n}:=\frac{1}{n}\sum_{i=1}^{n}1_{\{g_{n-1}(X_{i},D_{n,i})\neq Y_{i}\}},$
where $D_{n,i}$ is obtained from $D_{n}$ by omitting the pair $(X_{i},Y_{i})$. The Devroye–Wagner theorem [13, Theorem 24.2] says that
$\mathbb{E}[(\widehat{L}_{n}-L_{n})^{2}]\leq\frac{1}{n}+6\,\mathbb{P}(g_{n}(X_{n+1},D_{n})\neq g_{n-1}(X_{n+1},D_{n-1})).$ (2.5)
A very similar bound can be obtained from our Theorem 1.1, as follows. In addition to $g_{n-1}$, we need a symmetric prediction rule $g_{n-2}$ for samples of size n2n-2. Let $A(D_{n}):=\{(x,y):y=g_{n}(x,D_{n})\}$, and define $A^{\prime}$ and $A^{\prime\prime}$ similarly. Then
$\delta^{\prime}$ $=\mathbb{E}[\mu(A(D_{n})\Delta A^{\prime}(D_{n-1}))]$
$=\mathbb{E}[\mathbb{P}(\{Y_{n+1}=g_{n}(X_{n+1},D_{n})\}\Delta\{Y_{n+1}=g_{n-1}(X_{n+1},D_{n-1})\}|D_{n},X_{n+1})]$
$=\mathbb{P}(\{Y_{n+1}=g_{n}(X_{n+1},D_{n})\}\Delta\{Y_{n+1}=g_{n-1}(X_{n+1},D_{n-1})\})$
$\leq\mathbb{P}(g_{n}(X_{n+1},D_{n})\neq g_{n-1}(X_{n+1},D_{n-1})).$
Similarly,
$\delta^{\prime\prime}\leq\mathbb{P}(g_{n-1}(X_{n+1},D_{n-1})\neq g_{n-2}(X_{n+1},D_{n-2})).$
Bounding $\theta$ by $\frac{1}{4}$, we get the bound
$\mathbb{E}[(\widehat{L}_{n}-L_{n})^{2}]$ $\leq\frac{1}{4n}+4\,\mathbb{P}(g_{n}(X_{n+1},D_{n})\neq g_{n-1}(X_{n+1},D_{n-1}))$
$\qquad+4\,\mathbb{P}(g_{n-1}(X_{n+1},D_{n-1})\neq g_{n-2}(X_{n+1},D_{n-2})).$
In practice, this would be essentially be the same as the bound (2.5) from the Devroye–Wagner theorem. Still, it would be interesting to understand why there are two terms in our bound, versus only one in (2.5). In particular, would it be possible to improve Theorem 1.1 to have only one error term, corresponding to the comparison of $n$ and n1n-1? It is not immediately clear, since Theorem 1.1 has wider coverage than the Devroye–Wagner theorem. We leave this question for future investigation.
2.8 Connection to algorithmic stability
Roughly speaking, algorithmic stability is the notion that the output of an algorithm, whose input is a set of i.i.d. observations, should not change very much under omitting any one of the observations. It is clear that this idea is pertinent for applications of Theorem 1.1. Indeed, Theorem 1.1 says that under a certain kind of algorithmic stability, the leave-one-out estimate of the size of a random set is a good estimate. The various applications we have worked out in this paper are all about showing that the respective algorithms are stable. The main difference with the prior literature is that the usual versions of algorithmic stability mainly look at point estimates, whereas we are looking at random sets. We refer to Chapter 6 in the forthcoming monograph of Angelopoulos, Barber, and Bates [1] for references to the algorithmic stability literature.
@@ Acknowledgements
We thank Louigi Addario-Berry, David Aldous, Eugenio Regazzini, Richard Stanley, and Sandy Zabell for their help. PD was supported by NSF grant DMS-1954042. SC was supported in part by NSF grants DMS-2113242 and DMS-2153654.